OpenStudy (anonymous):

Can someone help please What is the area of a triangle with a base of 3.2 cm and a height of 9.6 cm? A. 6.4 cm B. 12.8 cm C. 15.36 cm D. 30.72 cm

Vocaloid (vocaloid):

area = (1/2)*base*height

OpenStudy (anonymous):

A = 1/2Bh. 3.2*9.6 = 30.72 30.72/2 = ?

OpenStudy (anonymous):

c.15.36

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

@nerdybird

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

im new so i dont know how too give medals